Hi,
I'm trying to solve a problem around coin flips. If the probability of a "win" is .48 and I'm not going to stop flipping until I've reached 10 wins or 10 losses, what are the odds of reaching each number?
Thanks in advance!!!
Dmitry

Answer by stanbon(75887)

I'm trying to solve a problem around coin flips. If the probability of a "win" is .48 and I'm not going to stop flipping until I've reached 10 wins or 10 losses, what are the odds of reaching each number?
----
odds for a win = [P(win)/P(not win)] = 0.48/0.52 = 12:13
===============
Cheers,
Stan H.